# Rydberg Atomic Receiver: Next Frontier of Wireless Communications

### v1

@@ III-D2 Instantaneous bandwidth
The communication rate of a RARE is restricted by several factors, such as the response time for a EIT process to become steady. As reported by exciting studies, the typical instantaneous bandwidth of a RARE is around 10 MHz [4], comparable to the bandwidth of 4G LTE (i.e., 1$\sim$20 MHz). Some recent efforts are endeavored to increase the RARE’s bandwidth. For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time [12]. This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6GHz systems specified by 5G standards (i.e., 20$\sim$100 MHz). To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
IV Recent Progress in RARE Techniques
@@
IV-A Frequency-division-Multiplexed RAREs
The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference and efficient utilization of spectrum resources. The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications. The state-of-the-art FDM-RAREs can be separated into two categories: standard and cross-band FDM-RAREs.
@@ IV-A1 Standard FDM-RARE
As illustrated in Fig. 5, a standard FDM-RARE aims to detect standard FDM signals, where all sub-carrier frequencies lie within a bandwidth of several MHz with a kHz-level frequency spacing. In this context, a uniform electron transition |$\ket{3}\leftrightarrow\ket{4}$ is sufficient to down-convert all sub-carrier signals into the Rabi frequency. The major challenge in facilitating standard FDM-RAREs is that the nonlinearity of Rabi frequency, $\Omega(t)$, w.r.t the EM wave destroys the orthogonality among different sub-carriers, resulting in significant intercarrier crosstalk. To address this issue, a deep learning model was trained in [10] to directly infer FDM symbols from the received probe-laser power. The experiment therein considered four subcarriers with a frequency spacing of 2 kHz in line-of-sight environment. It was shown that the learning approach can demodulate binary phase-shift keying symbols with over 99% accuracy.
IV-A2 Cross-band FDM-RARE
The cross-band FDM-RARE leverages the broad operating bandwidth of RAREs to serve as an integrated full-frequency communication platform. On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig. 5. To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$. The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire cross-band RF signal to the baseband. Utilizing this principle, researchers have demonstrated the co-detection of cross-band signals over 5 separate bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [6].
Figure 5: Standard and cross-band FDM-RAREs.
@@ IV-B Atomic-MIMO Communications
As a core technology in 5G and Beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains. This motivates the research on atomic-MIMO communications, which utilizes an array of RAREs for receiving wireless data to boost both the sensitivity and channel capacity of RAREs.
In the literature, a simpler atomic single-input-multiple-output (SIMO) system was developed in [13] for the first time to explore the diversity gain, wherein four RAREs were used to jointly recover AM symbols from a single-antenna user. Experiments validated that the receive SNR scales linearly with the number of RAREs, as in traditional systems. To further reap the multiplexing gain, the current authors have investigated the multi-user signal detection capability in atomic-MIMO systems [8], where the EIT-heterodyne scheme was adopted to retrieve Quadrature-Amplitude-Modulated (QAM) symbols. It was proved that atomic-MIMO exhibits a nonlinear phase-retrieval transmission model, based on which an expectation-maximization algorithm was designed to accurately recover multi-user signals, demonstrating the feasibility of atomic-MIMO communications. In a followup work, the optimal MIMO precoding has been designed to achieve the channel capacity of atomic-MIMO [14]. It made the assumption to the EIT-heterodyne scheme that the reference signal shares the same frequency with the data signal and is much stronger. It was shown that the optimal precoder should process the I/Q baseband symbols independently, and the channel capacity scales as $\min(N_{r}/2,N_{t})\log{\rm SNR}$ ) for $N_{r}$ atomic receive antennas and $N_{t}$ classic transmit antennas. The discount factor $\frac{1}{2}$ imposed on $N_{r}$ arises because the Rabi frequency, $\Omega(t)$, only captures the magnitude of EM waves.
IV-C RARE-empowered Wireless Sensing
Given that a RARE is capable of detecting extremely weak signals, it can be deployed in wireless sensing systems to refine the sensing accuracy by orders of magnitude. Specifically, one existing prototype senses a target’s vibration from both WiFi (2.4 GHz) and mmWave (28 GHz) signals [5]. Therein, the EIT-heterodyne scheme, as illustrated in Fig. 2, was used to recover the phase of an EM wave, ${\mathbf{E}}(t)$, and then retrieve the vibration information. Experiment results show that the RARE can increase the sensing accuracy by more than 10 times w.r.t conventional receivers implemented on the USRP, LimdSDR, and Intel 5300 WiFi card. In addition to vibration sensing, RARE can also be used to sense the angle-of-arrival (AoA) for the purpose of localization. By measuring the signal phases at two different locations via the EIT-heterodyne scheme, the AoA of the signal can be obtained with a precision of $2.1^{\circ}$ [4].
@@ IV-D Quantum Many-body RAREs for External-noise Mitigation
One intriguing approach leverages the quantum many-body effect of Rydberg atoms to counterintuitively convert the power of external noise into that of wireless signal, achieving a noise-enhanced RARE [15]. In the approach, a collection of Rydberg atoms are interacted with each other. This many-body interaction transforms the original linear Schrödinger equation into a nonlinear one, which is shown to exhibit a bistability behavior. The incoming wireless signal could stimulate the system to switch between the two stable states. Then, by introducing a certain amount of external noise into the receiver, the switching magnitude can be amplified such that the output SNR is drastically improved. Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise. This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the nonlinear many-body interactions in Rydberg atoms.
V Future Directions
@@
Despite the rapid development of RARE in recent years, its integration with practical communication systems remains at a nascent stage. Numerous wireless-communication related open problems should be addressed to unleash the full potential of RAREs. A few are discussed as follows.
V-A Generalized FDM-RARE Techniques
While initial solutions [10, 6] exist for both standard and cross-band FDM-RAREs, the systematic design of FDM-RARE technology, involving channel estimation, signal detection, and synchronization, remains an unaddressed and crucial challenge. For standard FDM-RAREs, generalizing the learning method in [10] to accommodate hundreds of subcarriers in a multi-path environment is difficult due to the unexplainability of neural networks. Regarding cross-band FDM-RAREs, it is noteworthy that the quantum system has evolved from a four-level system to an $N$-level system [6]. However, at the time of writing, a comprehensive transmission model for this new $N$-level system has yet to be developed, and the potential crosstalk between cross-band signals remains unexplored. To address the above issues, it is essential to start from first principles and systematically develop transmission methods tailored for FDM-RAREs by thoroughly investigating their nonlinear transmission models.
V-B Enhanced Atomic-MIMO Communications
Figure 6: Enhanced atomic-MIMO communications.
The research on atomic-MIMO communication is still at its early stage. Numerous signal processing techniques at both the transmitter (Tx) and receiver (Rx) sides can be investigated to improve the communication performance, as shown in Fig. 6.
First, consider the Rx side. Recall that a RARE might be vulnerable to external interference. Apart from the aforementioned quantum many-body method, the development of signal combining algorithms tailored for atomic-MIMO presents a viable approach to mitigating inter-user interference. Furthermore, using RAREs for precise estimation of far-field and near-field channels is crucial for effective atomic-MIMO communications, but remains an uncharted area. Lastly, investigating the integration of atomic-MIMO and FDM-RARE techniques could unlock RARE’s capability of full multi-user and multi-frequency signal detection.
Next, developing Tx-side signal processing methods optimized for RAREs can also significantly improve the communication performance. To this end, one could relax some restrictive assumptions made in the MIMO precoding study in [14] we discussed earlier and explore the general EIT-heterodyne scheme, where the reference signal can have a different frequency and comparable strength to the user signal. Within this framework, optimizing the joint power allocation among the unknown user signal, ${\mathbf{E}}(t)$, and known reference signal, ${\mathbf{E}}_{r}(t)$, designing precoding for ${\mathbf{E}}(t)$, and refining the waveform of ${\mathbf{E}}_{r}(t)$, are all potential approaches for capacity enhancement. Moreover, it is promising to introduce diverse beamforming technologies to cooperate with RAREs, such as fluid-antenna, reconfigurable-intelligent-surface, and continuous-aperture-array empowered beamforming.
V-C Atomic Integrated Communication and Sensing
The extreme sensitivity of RAREs to both communication and sensing signals can facilitate the exciting application of atomic Integrated Communication and Sensing (ISAC). To advance this direction, the signal waveform utilized in conventional ISAC systems, like the orthogonal-chirp-division-multiplexing signal, need to be adjusted to align with the nonlinear transmission model revealed by Rabi frequency. Additionally, it is promising to exploit the cross-band response capability of Rydberg atoms in atomic-ISAC systems. By exciting electrons to proper energy levels, it becomes feasible to simultaneously utilize sub-6GHz signals to ensure communication coverage and high-frequency signals to refine sensing accuracy.
V-D Advancing Quantum Many-body RAREs
A more thorough exploration of the quantum many-body effects of Rydberg atoms is essential for mitigating external noises. The preliminary study in [15] focused on periodic sinusoidal signals and white external noise, with the evaluation metric limited to SNR. Holistic theories are required to investigate the characteristics of this quantum many-body RARE under the influence of aperiodic communication signals (such as QAM signals and FDM signals), and thereby quantitatively analyze the reduction of practical color external noise (such as inter-user interference). In addition, analysis of the input-output mutual information should be carried out to unveil the communication gain brought by the many-body effect, rather than solely relying on the SNR. Last, it is intractable to derive an analytical transmission model from the nonlinear Schrödinger equation. Thereafter, it might be necessary to use artificial intelligence technologies, such as physics-informed neural networks, to develop signal detection and channel estimation algorithms for quantum many-body RAREs.
VI Conclusions
@@
This article provides an overview of the fundamentals, advantages, progresses, and future directions for RARE-aided wireless communications. Representing a paradigm shift in receiver architecture, RAREs detect classical EM signals via quantum phenomena, contributing to improved sensitivity, enhanced coverage, and broadened operating bandwidth. RAREs can seamlessly integrate with both classical techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance. However, the advancement of RAREs requires more efforts from the communication community to systematically develop mature transmission models, theoretical analysis frameworks, and efficient communications algorithms, making RARE a practically useful technology for future wireless networks.
@@
Mingyao Cui is a PhD student at The University of Hong Kong, Hong Kong.

### v2

@@ III-C2 Instantaneous bandwidth
Restricted by the long response time for an EIT process to become steady, the instantaneous bandwidth of a RARE is typically narrower than 10 MHz [7]. This bandwidth is comparable to that of 4G LTE systems (1$\sim$20 MHz). Recent efforts are endeavored to increase the RARE’s bandwidth. For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time (see [139] of [2]). This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz). To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
IV State-of-the-art RARE Techniques
@@
IV-A Frequency-division Multiplexed RAREs
The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference. The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications. The state-of-the-art FDM-RAREs can be separated into two categories: standard and multi-band FDM-RAREs.
@@ IV-A1 Standard FDM-RARE
As illustrated in Fig. 4, a standard FDM-RARE aims to detect standard FDM signals, where all sub-carrier frequencies are compacted within a bandwidth of several MHz. In this context, a uniform electron transition |$\ket{3}\leftrightarrow\ket{4}$ is sufficient to down-convert all sub-carrier signals into the Rabi frequency of quantum state. The major challenge in facilitating standard FDM-RAREs is that the nonlinearity of Rabi frequency, $\Omega(t)$, w.r.t the EM wave destroys the orthogonality among different sub-carriers, resulting in significant intercarrier crosstalk. To address this issue, a deep learning model was trained in [10] to directly infer FDM symbols from the received probe-laser power. The experiment therein considered four subcarriers with a frequency spacing of 2 kHz in line-of-sight environment. It was shown that the learning approach can demodulate binary phase-shift keying symbols with over 99% accuracy.
IV-A2 Multiband FDM-RARE
The multiband FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform. On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig. 4. To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$. The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire multi-band RF signal to the baseband. Utilizing this principle, researchers have demonstrated the co-detection of 5 separate frequency bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [5].
Figure 4: Standard and multiband FDM-RAREs.
@@ IV-B Atomic-MIMO Communications
As a core technology in 5G and beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains. As a representative, the recent study in [12] has demonstrated a good performance with an 20 Gbps 1500 m transmission distance in free-space-optical communications using advanced MIMO diversity combining codes. The advancement of MIMO thus motivates the research on atomic-MIMO communications. Its key idea is to employ an array of RAREs, in place of traditional antenna arrays, for signal detection. In this context, both the high sensitivity of Rydberg atoms and the high spectral efficiency of MIMO can be leveraged.
In the literature, a simpler atomic single-input-multiple-output (SIMO) system was developed in [13] for the first time to explore the diversity gain, wherein four RAREs were used to jointly recover AM symbols from a single-antenna user. Experiments validated that the receive SNR scales linearly with the number of RAREs, as in traditional systems. To further reap the multiplexing gain, [8] has investigated the multi-user signal detection capability of atomic-MIMO. It was proved that atomic-MIMO exhibits a nonlinear phase-retrieval transmission model. The nonlinear characteristic negates the efficacy of traditional zero-forcing and minimum-mean-square-error linear detectors, posing a challenge to atomic-MIMO signal detection. To this end, an Expectation-Maximization Gerchberg-Saxton algorithm was proposed to iteratively retrieve multi-user signals with a low complexity, demonstrating the feasibility of atomic-MIMO.
IV-C RARE-empowered Quantum Wireless Sensing
Given that a RARE is capable of detecting extremely weak signals, it can be deployed in wireless sensing systems to refine the sensing accuracy by orders of magnitude. Specifically, one existing prototype senses a target’s vibration from both WiFi (2.4 GHz) and mmWave (28 GHz) signals [4]. Therein, the heterodyne sensing scheme, as illustrated in Fig. 2, was used to recover the phase of an EM wave and then retrieve the subtle-vibration of target. Experiment results show that the RARE can increase the sensing accuracy by more than 10 times as compared with conventional receivers implemented on the USRP, LimdSDR, and Intel 5300 WiFi card.
In addition, RARE can also be utilized to sense angle-of-arrivals (AoA) and time-of-arrivals (ToA) for the purpose of localization. A self-heterodyne sensing technique was introduced in [14] to advance the current heterodyne sensing for high-resolution localization. This technique leverages the self-interference produced by the transmitted sensing signal as the reference signal, ensuring their waveforms to be identical. By doing so, it effectively overcomes the narrow instantaneous bandwidth limitation of RARE, enhancing the sensing bandwidth up to 100 MHz. Experiments show that a self-heterodyne RARE reduces the localization error by 20 dB compared with classic receivers.
@@ IV-D Quantum Many-body RAREs for External-noise Mitigation
One intriguing approach leverages the quantum many-body effect of Rydberg atoms to counterintuitively convert the power of external noise into that of wireless signal, achieving a noise-enhanced RARE [15]. In the approach, a collection of Rydberg atoms are interacted with each other. This many-body interaction transforms the original linear Schrödinger equation into a nonlinear one, which is shown to exhibit a bistability behavior. The incoming wireless signal could stimulate the system to switch between the two stable states. Then, by introducing a certain amount of external noise into the receiver, the switching magnitude can be amplified such that the output SNR is drastically improved. Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise. This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
V RARE for Multiband Sensing and Communication (RARE-MSAC)
Figure 5: A single RARE is used to simultaneously detect communication signals at low and medium frequency bands, and sensing signals at high frequency bands.
We leverage the advancement of RARE to introduce a novel application in PHY wireless systems——the RARE-MSAC. The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multiband signals with diverse sensing and communication (S&C) requirements. As illustrated in Fig. 5, a single RARE is employed. The atoms are prepared in an appropriate initial Rydberg state, |$\ket{3}$, where the transitions from |$\ket{3}$ to other Rydberg states, |$\ket{n},\forall n\geq 4$, can be coupled to distinct frequency bands. In this context, the RARE is capable of concurrently detecting communication signals at low ($<6{\rm GHz}$) or medium bands ($7{\rm GHz}-24{\rm GHz}$) to ensure coverage. Simultaneously, it can also capture sensing signals operating in the mmWave and THz bands to perform sensing tasks with varying granularity requirements, ranging from ToA estimation to heartbeat detection. It is amazing that all these S&C services are achieved by a single RARE device, which is nearly impossible for classic receivers due to the largely separated frequencies.
Fig. 6 compares the MSAC performance of the RARE and classic receivers. We consider a dual-band S&C system. A sub-6G device transmits communication signals operating at $3.213\>{\rm GHz}$ with an instantaneous bandwidth of $500\>{\rm kHz}$. Simultaneously, a mmWave device transmits sensing signals operating at $30.618\>{\rm GHz}$ with a bandwidth of $100\>{\rm kHz}$. Similar to [4], the sensing signal is reflected by a target for detecting its vibration . A single RARE is deployed to detect these dual-band signals for providing S&C services. The Rydberg states are configured as |$\ket{3}=60D_{5/2}$, |$\ket{4}=61P_{3/2}$, and |$\ket{5}=62P_{3/2}$, respectively. The electron transition $60D_{5/2}\rightarrow 61P_{3/2}$ has a dipole moment of $2.04\times 10^{-26}\>{\rm C\cdot m}$ and resonates with the $3.213\>{\rm GHz}$ communication signal, while the transition $60D_{5/2}\rightarrow 62P_{3/2}$ has a dipole moment of $6.24\times 10^{-27}\>{\rm C\cdot m}$ and resonates with the $30.618\>{\rm GHz}$ sensing part. The power of the probe laser and coupling laser are set as $120{\rm\mu W}$ and $34\>{\rm mW}$, respectively. For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the sub-6G communication service and “CR2” for the mmWave sensing service in Fig. 6. The transmission power for the two signals increases from $10\>{\rm dBm}$ to $40\>{\rm dBm}$. As observed from the experimental results of Fig. 6, the spectrum efficiency in the communication service achieved by RARE is $\sim 6{\rm bps/Hz}$ higher than that of the classic receiver under Rayleigh fading channels. Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by the classic receiver by two orders of magnitude. Notably, these performance gains are accomplished by a single RARE without requiring additional RF circuits.
Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers, denoted as “CR 1” and “CR 2” respectively.
VI Future Research Directions
@@
Despite the rapid development of RARE, its integration with practical communication systems remains at a nascent stage. Numerous physical-layer research problems deserve in-depth investigation to unleash the full potential of RAREs.
  • More accurate transmission model. The current signal transmission model of RARE mainly considers the interaction between Rydberg atoms and EM waves. In practice, the Rydberg-Rydberg interaction can significantly influence the dynamics of quantum states, which should be thoroughly studied to make the signal transmission model of RARE more accurate.
  • Joint transmitter and RARE optimization. As RARE changes the signal transmission model from linear to nonlinear, traditional signal processing methods become suboptimal. It is essential to develop new transmission techniques tailored for both the RARE and transmitter, including modulation, precoding, channel estimation, synchronization, interference mitigation, and so on.
  • Physical layer security. RARE also holds promise in enhancing physical layer security. Taking advantage of the abundant energy levels, a RARE can flexibly switch the frequency band it is listening to. This capability enables the implementation of high-speed and ultra-broadband frequency hopping communication, which can effectively protect the privacy of wireless data.
  • Satellite communication and low-altitude economy. Another promising application of RARE is to assist satellite communication and low-altitude economy. The high sensitivity of RARE not only enables the expansion of satellite signal coverage but also demonstrates the potential to enhance the precision of sensing and navigation for unmanned aerial vehicles.
VII Conclusions
@@
This article provides an overview of RARE-aided wireless communications. Representing a paradigm shift in receiver architecture, RAREs detect EM signals via quantum phenomena, contributing to improved sensitivity and broadened detecable frequency range. RAREs can seamlessly integrate with both classic techniques, such as FDM and MIMO, as well as quantum techniques, like the many-body effect, to enhance communication performance. The potential of RARE in multiband S&C systems is revealed through experimental results. Finally, we have highlighted several promising future research directions on RARE.
@@
Mingyao Cui is a PhD candidate at The University of Hong Kong, Hong Kong.